OpenStudy (anonymous):

What are the possible remaining angle measures in triangle ABC with A=40, AC=60, and BC=45? a.

OpenStudy (anonymous):

We can use the law of sines for this one. We have a side length and two angles. If we get back a sensible answer i.e. >0 and real valued then we should be good to go. The law of sines is \[\frac{\sin(a)}{\alpha}=\frac{\sin(b)}{\beta}=\frac{\sin(c)}{\gamma} \]